• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name aniol

    Log entry time 00:02:32 on June17,2004

    Entry number 126019

    keyword=shift summary

    shift starts with no septa and no beam
    Looking at the end of run print outs it appears that the last good production run ended with run 2189 at 8:32 this morning
    16:18 hall in power permit
    17:00 moller measurements start
    20:00 start compton tune
    the left septum is still in manual fill mode and is still reading ~4.3% fill with Helium - hence it is off. The right septum is working.
    20:08 start r2196 production- compton tuning in progress
    20:27 end run r2196
    20:30 - called Hall B  about dithering. They want to wait 1/2 hour.
    20:37 start run r2197 prod run with RHRS, compton tuning in progress
    20:50 looks like a very low background in the compton tune,<100Hz
    21:04 end run 2197 to do dithering
    21:06 start dithering run 2198, dithering working properly after evt 22000 - note additional hand written instruction on procedure sheet
    21:30 end dithering run 2198
    21:32 start prod run 2199
    22:28 end run 2199, 100K
    22:36 start run 2200, short run to check new adc
    Bryan replaces a bad adc (reads out bpm10 )
    22:47 end run 2200
    22:51 start run 2201 prod run
    23:45 end run 2201 98.5K - error message appears on happex feedback monitor
    kill coda
    23:49 start run 2202 production run on RHRS
    
    
    
    Coulomb*spec score  total
    r2196 0.036C*1spec   0.036
    r2197 0.046*1spec    0.082
    r2199 0.102*1spec    0.184
    r2201 0.092*1spec    0.276